OpenStudy (anonymous):

solve (x+10)(x-17)(x+3)>=0 what is the solution set or is the solution all real numbers

OpenStudy (anonymous):

cubic polynomial with zeros at -10, -3 and 17

OpenStudy (sriram):

[-10,-3] U [17,INFINITY)

OpenStudy (anonymous):

is it an inequity all real numbers or no solution?

OpenStudy (anonymous):

so it will be negative on \[(-\infty,-10)\] positive on \[(-10,-3)\] negative on \[(-3,17)\] and positive on \[(17,\infty)\]

OpenStudy (anonymous):

solution is what sriram said

OpenStudy (anonymous):

which of the three options i gave is the solution and if its a complex inequity how would it be written?

OpenStudy (id21):

As you can see from attachment. The solution where graphic is >=0 is \[[-10,-3]\cup[17,+\infty)\]

OpenStudy (anonymous):

i sense from your questions that you may not be clear on what this means. it is clear what the question is asking?